OpenStudy (anonymous):

Help Please! find the 2nd term of an arithmetic sequence if t1=0 amd t5=16?

sam (.sam.):

Well arithmetic formula is \[T_n=a_1+(n-1)d\] -------------------------- \[T_1=a_1+(1-1)d=0 \] getting \(a_1=0\) -------------------------- \[T_5=0+(5-1)d=16\] solve for d then apply arithmetic formula again for t2

OpenStudy (anonymous):

4?

sam (.sam.):

yes now use \[T_n=a_1+(n-1)d\] for n=2